What is the integral of 1/(z-1) (z-2) (2-4) dz where C is |z| = 3?


1
Expert's answer
2022-02-10T13:02:34-0500

"\displaystyle
\int_{|z|=3}\frac{1}{(z-1) (z-2) (z-4)}\ dz"

"\displaystyle
=\frac{1}{3}\int_{|z|=3}\frac{1}{z-1}\ dz+\frac{1}{6}\int_{|z|=3}\frac{1}{z-4}\ dz-\frac{1}{2}\int_{|z|=3}\frac{1}{z-2}\ dz"

But,

"\displaystyle
\int_{|z|=3}\frac{1}{z-1}\ dz=2\pi i", by Cauchy integral theorem.


"\displaystyle
\int_{|z|=3}\frac{1}{z-4}\ dz=0", by Cauchy-Goursat theorem.


"\displaystyle
\int_{|z|=3}\frac{1}{z-2}\ dz=2\pi i", by Cauchy integral theorem.


Thus,

"\displaystyle
\int_{|z|=3}\frac{1}{(z-1) (z-2) (z-4)}\ dz"

"\displaystyle
=\frac{1}{3}\int_{|z|=3}\frac{1}{z-1}\ dz+\frac{1}{6}\int_{|z|=3}\frac{1}{z-4}\ dz-\frac{1}{2}\int_{|z|=3}\frac{1}{z-2}\ dz"

"\displaystyle
=\frac{1}{3}\times 2\pi i+\frac{1}{6}\times 0-\frac{1}{2}\times2\pi i=-\frac{\pi i}{3}"